write an equation in slope intercept form for th line that satisfies the following condition. passes through (10,16) perpendicular to the graph of 9x+12y=15
Answer by mananth(16949)

write an equation in slope intercept form for th line that satisfies the following condition. passes through (10,16) perpendicular to the graph of 9x+12y=15
..
9x+12y=15
12y=-9x+15
/12
y=-9/12 x +15/9
y = -3/4 *x +5/3
...
slope = -3/4
a line perpendicular to the given line will be 4/3
..
point( 10,16)
(y-16)=4/3(x-10)
y-16 = 4/3 *x -40/3
y= 4/3*x-40/3+16
y=4x/3 +8/3
..

..
m.ananth@hotmail.ca